OpenStudy (anonymous):

How do I do e^2x-7e^x+10=0

OpenStudy (anonymous):

how do you solve it for x you mean? it's a quadratic in e^x e^x = U U^2 -7U +10 =0 solve for U solve" U= e^x" for x

OpenStudy (anonymous):

What?.

OpenStudy (anonymous):

(e^x)^2 -7(e^x) +10 =0

OpenStudy (anonymous):

agree with that?

OpenStudy (anonymous):

from the property of exponents: (x^a)^b = x^(a*b)